# High Probability Bound for Cross-Learning Contextual Bandits with Unknown Context Distributions

### v1

@@ 5       Play $A_{t}\sim s_{1,c_{t}}$
6       $\hat{f}_{2}\leftarrow\hat{f}_{2}+\frac{s_{2,c_{t}}}{2L}$
@@ 8       $\hat{f}_{e+1}\leftarrow 0$
10             Set $p_{t,c}=\operatorname*{arg\,min}_{x\in\Delta([K])}\left(\left\langle x,\sum_{s=1}^{t-1}\hat{\ell}_{s}(c)\right\rangle-\eta^{-1}F(x)\right)$
@@ 19            $t_{f},t_{\ell}\leftarrow\mathsf{RandPerm}(t,t+1)$
20             $\hat{f}_{e+1}\leftarrow\hat{f}_{e+1}+\frac{s_{e+1,c_{t_{f}}}}{2(L/2)}$
@@
22             Set ^$\hat{\ell}_{t_{\ell},c}(a)=\frac{2\ell_{t_{\ell},c}(a)}{\hat{f}_{e}(a)+\frac{3}{2}\gamma}\mathbb{I}\left(A_{t}=a,S_{t}=1\right)$
@@ 23      $s_{e+2}\leftarrow p_{t}$
3.2 Overview of Our Analysis
In this subsection, we briefly overview our proof of Theorem 1. The full proof can be found in the appendix.
@@ 3.2 Overview of Our Analysis
$+$ $\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\langle p_{t,c}-\pi_{c},\hat{\ell}_{t,c}\rangle}_{\textbf{ftrl}}+\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\tilde{\ell}_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{3}}$
$+$ $\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\tilde{\ell}_{t,c}-\hat{\ell}_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{4}}+\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\ell_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{5}}.$
Our decomposition is different from the decomposition in Schneider and Zimmert (2023), this difference is essential for deriving a high-probability bound. The key difference lies in the term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\tilde{\ell}_{t,c}\right\rangle$ in the $\operatorname{\mathbf{bias}}_{2}$ term of Schneider and Zimmert (2023), which becomes $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\ell_{t,c}\right\rangle$ in the $\operatorname{\mathbf{bias}}_{5}$ term here. This difference saves a $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\tilde{\ell}_{t,c}-\ell_{t,c}\right\rangle$ term, which is crucial for deriving a high-probability bound.
In our decomposition, the $\operatorname{\mathbf{bias}}_{1}$ term refers to the bias introduced by replacing the regret over the entire time horizon with that over $\operatorname{\mathcal{T}^{\ell}}$, and the $\operatorname{\mathbf{bias}}_{2}$ term refers to the bias introduced by replacing regret with its linearization. Both of these terms are not hard to bound using standard concentration inequalities. Furthermore, the ftrl and $\operatorname{\mathbf{bias}}_{3}$ terms are standard in the analysis of high-probability bounds for bandit algorithms. These two terms are not hard to bound using techniqes from EXP3-IX (Neu, 2015) and Schneider and Zimmert (2023).
The terms of interest to bound are $\operatorname{\mathbf{bias}}_{4}$ and $\operatorname{\mathbf{bias}}_{5}$. These two terms can be bounded using similar methods. Here we give a sketch of the analysis of the $\operatorname{\mathbf{bias}}_{5}$ term. Details can be found in the appendix.
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\hat{\ell}_{t,c}-\ell_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\ell_{t,c}(\pi_{c})\right)$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right).$
In this decomposition, the two terms correspond to different components of the bias of the estimator ^$\hat{\ell}_{t,c}$. The first term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$ corresponds to the bias introduced by constructing the importance estimator $\hat{f}_{e}(a)$. The second term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\hat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$ corresponds to the bias introduced from the randomness in sampling $a_{t}$ from $q_{t}$. Once again, the analyses of these two terms follow the same principle. We briefly introduce the analysis of the first term for the sake of simplicity. Details can be found in the appendix.
To bound the first term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$, we will use the key observation mentioned at the beginning of this section: different epochs in Algorithm 1 are only weakly dependent on each other. To use this observation rigorously, we introduce an important technical tool. With a slight abuse of notation, we define a filtration $\left\{\operatorname{\mathcal{H}}_{e}\right\}_{e}$, in which for each epoch $e$, the $\sigma$-algebra $\operatorname{\mathcal{H}}_{e}$ is generated by all randomness in epochs $11,\dots,e-$ 1 and the randomness in $\operatorname{\mathcal{T}_{e}^{\ell}}$. That is, the $\sigma$-algebra $\operatorname{\mathcal{H}}_{e}$ is generated precisely by the context $c_{t}$, the random seed used in sampling $a_{t}\sim q_{t,c_{t}}$, and the random seed used in sampling $S_{t}\sim\mathcal{B}\left(\frac{s_{e,c_{t}}(a_{t})}{2q_{t,c_{t}}(a_{t})}\right)$ for $t\leq(e-1)L$ and $t\in\operatorname{\mathcal{T}_{e}^{\ell}}$. Note that for each epoch $e$, the $\sigma$-algebra $\operatorname{\mathcal{H}}_{e}$ excludes the randomness in $\operatorname{\mathcal{T}_{e}^{f}}$. This exclusion is crucial for characterizing the weak dependence structure between epochs.
@@
Then the first term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$ is exactly $\sum_{e}\operatorname{\mathrm{Bias5}}_{e}$. Our key observation is that, not only
$\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\hat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)\right]\leq 0$
@@ Appendix A Useful Lemmas
$L_{e}\triangleq\operatorname{\mathbbm{1}}\left(\max_{c\in[C],a\in[K]}\sum_{t\in\mathcal{T}_{e}}\tilde{\ell}_{t,c}(a)\leq L+\frac{\iota}{\gamma}\right).$
@@
$\frac{1}{2}\leq\frac{f_{e}(a)+\gamma}{\hat{f}_{e}(a)+\frac{3}{2}\gamma}\leq 2.$
@@
$\widetilde{p}_{t,c}\triangleq\operatorname*{arg\,min}_{p\in\Delta([K])}\left\langle p,\sum_{e^{\prime}=1}^{e-1}\sum_{s\in\mathcal{T}_{e^{\prime}}}\widehat{\ell}_{sc}+\sum_{t^{\prime}\in\mathcal{T}_{e},t^{\prime}<t}\tilde{\ell}_{t^{\prime}c}\right\rangle-\eta^{-1}F(p)$
@@
It is easy to see that $\widetilde{p}_{t,c}\propto s_{e+1,c}\circ\exp\left(-\eta\sum_{t^{\prime}\in\mathcal{T}_{e},t^{\prime}<t}\tilde{\ell}_{t^{\prime}c}\right).$
@@
$2s_{e,c}(a)\geq p_{t,c}(a)\geq s_{e,c}(a)/2\qquad\text{and}\qquad 2s_{e,c}(a)\geq\tilde{p}_{t,c}(a)\geq s_{e,c}(a)/2\,.$
@@
$\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]\leq 4f_{e}(a)\qquad\text{and}\qquad\mathbb{E}_{c\sim\nu}[\tilde{p}_{t,c}(a)]\leq 4f_{e}(a)\,.$
@@
We define $p_{t}(a)\triangleq\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ and $\tilde{p}_{t}(a)\triangleq\mathbb{E}_{c\sim\nu}[\tilde{p}_{t,c}(a)]$ for each time step $t$ and each arm $a$.
@@
$-\frac{\gamma}{f_{e}(a)}\leq\mathbb{E}\left[\frac{f_{e}\left(a\right)-\hat{f}_{e}\left(a\right)-\frac{1}{2}\gamma}{\hat{f}_{e}\left(a\right)+\frac{3}{2}\gamma}F_{e}\Big{|}\operatorname{\mathcal{H}}_{e-1}\right]\leq 0.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c}-\widetilde{p}_{t,c},\tilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\right\rangle G\leq\frac{98KT\iota}{L}+\frac{\gamma^{2}LKT}{\iota}.$
@@
Decomposing all time steps into consecutive pairs, specifically, decomposing $\{1,2,\dots,T\}$ into $\{(1,2),(3,4),(5,6),\dots,(t-1,t),\dots,(T-1,T)\}$. Constructing a surrogate loss sequence {~$\{\tilde{\ell}_{s}\}_{s=1}^{\frac{T}{2}}$ such that for each surrogate time step $s$ the loss vector ~$\tilde{\ell}_{s}$ is uniformly sampled from the pair of true loss vector $(\ell_{2s-1},\ell_{2s})$. Denote the time step sampled from the pair $(2s-1,2s)$ as $s_{\ell}$. For any constant $\delta\in(0,1)$ and any bandit algorithm such that in each pair of time steps $(t-1,t)$, the algorithm takes actions $a_{t-1}$ and $a_{t}$ from the same distribution $p_{t-1}=p_{t}$, we have
$\sum_{t=1}^{T}\left(\ell_{t,c_{t}}(a_{t})-\ell_{t,c_{t}}(\pi_{c_{t}})\right)\leq 2\sum_{s=1}^{\frac{T}{2}}\left(\tilde{\ell}_{s,c_{s_{\ell}}}(a_{s_{\ell}})-\tilde{\ell}_{s,c_{s_{\ell}}}(\pi_{c_{s_{\ell}}})\right)+2\sqrt{T\log(\frac{1}{\delta})}$
@@
$-2\left(\tilde{\ell}_{s,c_{s_{\ell}}}(a_{s_{\ell}})-\tilde{\ell}_{s,c_{s_{\ell}}}(\pi_{c_{s_{\ell}}})\right).$
Consider the filtration $\{\tilde{H}_{s}\}_{s=1}^{\frac{T}{2}}$ such that for each $s$ the $\sigma$-field $\tilde{H}_{s}$ is generated by the randomness within $c_{t}$ and $a_{t}$ for $t\leq 2s$ and the randomness within sampling from pair $(2\tau-1,2\tau)$ for each $\tau\leq s$. It is easy to see that the sequence$\{Y_{s}\}_{s=1}^{\frac{T}{2}}$ forms a martingale difference sequence adapted to the filtration $\{\tilde{H}_{s}\}_{s=1}^{\frac{T}{2}}$. Moreover, it is also easy to see that $|Y_{s}|\leq 2$. Using Azuma-Hoeffding’s inequality, for any constant $\delta\in(0,1)$, we have

### v2

@@ 5       Play $A_{t}\sim s_{1,c_{t}}$
6       $\widehat{f}_{2}\leftarrow\widehat{f}_{2}+\frac{s_{2,c_{t}}}{2L}$
@@ 8       $\widehat{f}_{e+1}\leftarrow 0$
10             Set $p_{t,c}=\operatorname*{arg\,min}_{x\in\Delta([K])}\left(\left\langle x,\sum_{s=1}^{t-1}\widehat{\ell}_{s}(c)\right\rangle-\eta^{-1}F(x)\right)$
@@ 19            $t_{f},t_{\ell}\leftarrow\mathsf{RandPerm}(t,t+1)$
20             $\widehat{f}_{e+1}\leftarrow\widehat{f}_{e+1}+\frac{s_{e+1,c_{t_{f}}}}{2(L/2)}$
@@
22             Set ^$\widehat{\ell}_{t_{\ell},c}(a)=\frac{2\ell_{t_{\ell},c}(a)}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}\mathbb{I}\left(A_{t}=a,S_{t}=1\right)$
@@ 23      $s_{e+2}\leftarrow p_{t}$
4 Main Result and Analysis
The main result of our paper is the following theorem.
Theorem 1 (Formal).
For any $\delta\in(0,1)$, Algorithm 1 with parameters choice $\iota=2\log(8KT\frac{1}{\delta})$, $L=\sqrt{\frac{\iota KT}{\log(K)}}=\widetilde{\Theta}(\sqrt{KT\log\frac{1}{\delta}})$, $\gamma=\frac{16\iota}{L}=\widetilde{\Theta}(\sqrt{\frac{\log(1/\delta)}{KT}})$, and $\eta=\frac{\gamma}{2(2L\gamma+\iota)}=\widetilde{\Theta}(1/\sqrt{KT\log(1/\delta)})$ yields a regret bound of
$\operatorname{\mathrm{Reg}}(\pi)=\widetilde{O}\left(\sqrt{KT\log\frac{1}{\delta}}\right)$
with probability at least $1-\delta$ for any policy $\pi$.
In what follows, we briefly overview our proof of Theorem 1. The full proof can be found in the appendix.
4.1 Regret Decomposition
@@
$+$ $\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\langle p_{t,c}-\pi_{c},\widehat{\ell}_{t,c}\rangle}_{\textbf{ftrl}}+\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\ell_{t,c}-\widetilde{\ell}_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{3}}$
$+$ $\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{4}}+\underbrace{2\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\widehat{\ell}_{t,c}-\ell_{t,c}\right\rangle}_{\operatorname{\mathbf{bias}}_{5}}.$
In our decomposition, the $\operatorname{\mathbf{bias}}_{1}$ term refers to the bias introduced by replacing the regret over the entire time horizon with that over $\operatorname{\mathcal{T}^{\ell}}$, and the $\operatorname{\mathbf{bias}}_{2}$ term refers to the bias introduced by replacing regret with its linearization. Both of these terms are not hard to bound using standard concentration inequalities. Furthermore, the ftrl and $\operatorname{\mathbf{bias}}_{3}$ terms are standard in the analysis of high-probability bounds for bandit algorithms. These two terms are not hard to bound using techniqes from EXP3-IX (Neu, 2015; Schneider and Zimmert, 2023). The $\operatorname{\mathbf{bias}}_{4}$ and $\operatorname{\mathbf{bias}}_{5}$ terms correspond to the bias introduced by constructing the importance estimator $\widehat{f}_{e}(a)$. These two terms are the terms of interest to bound.
Our decomposition is different from the decomposition in Schneider and Zimmert (2023). This difference is essential for deriving a high-probability bound. The key difference lies in the $\operatorname{\mathbf{bias}}_{5}$ term here. This term saves a $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\widetilde{\ell}_{t,c}-\ell_{t,c}\right\rangle$ term from the $\operatorname{\mathbf{bias}}_{2}$ term in the decomposition given by Schneider and Zimmert (2023), which is crucial for deriving a high-probability bound.
4.2 Identifying a Prototypical Term
The terms of interest to bound are $\operatorname{\mathbf{bias}}_{4}$ and $\operatorname{\mathbf{bias}}_{5}$. These two terms can be bounded using similar methods. Here we take the $\operatorname{\mathbf{bias}}_{5}$ term as a prototypical term and give a sketch of its analysis. Details can be found in the appendix.
@@ 4.2 Identifying a Prototypical Term
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle\pi_{c},\widehat{\ell}_{t,c}-\ell_{t,c}\right\rangle$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\ell_{t,c}(\pi_{c})\right)$
$=$ $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$
$+\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right).$
In this decomposition, the two terms correspond to different components of the bias of the estimator ^$\widehat{\ell}_{t,c}$. The first term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$ corresponds to the bias introduced by constructing the importance estimator $\widehat{f}_{e}(a)$. The second term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\widehat{\ell}_{t,c}(\pi_{c})-\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]\right)$ corresponds to the bias introduced from the randomness in sampling $a_{t}$ from $q_{t}$. Once again, the analyses of these two terms follow the same principle. We take the first term
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$
as a prototypical term and give a sketch of its analysis for the sake of simplicity. Details can be found in the appendix.
4.3 Analysis of the Prototypical Term
To bound the term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$, we will use the key observation mentioned at Section 1.1: different epochs in Algorithm 1 are only weakly dependent on each other. To use this observation rigorously, we introduce an important technical tool. With a slight abuse of notation, we define a filtration $\left\{\operatorname{\mathcal{H}}_{e}\right\}_{e}$, in which for each epoch $e$, the $\sigma$-algebra $\operatorname{\mathcal{H}}_{e}$ is generated by all randomness in epochs $11,\dots,e-$ 1 and the randomness in $\operatorname{\mathcal{T}_{e}^{\ell}}$. That is, the $\sigma$-algebra $\operatorname{\mathcal{H}}_{e}$ is generated precisely by the context $c_{t}$, the random seed used in sampling $a_{t}\sim q_{t,c_{t}}$, and the random seed used in sampling $S_{t}\sim\mathcal{B}\left(\frac{s_{e,c_{t}}(a_{t})}{2q_{t,c_{t}}(a_{t})}\right)$ for $t\leq(e-1)L$ and $t\in\operatorname{\mathcal{T}_{e}^{\ell}}$. Note that for each epoch $e$, the $\sigma$-algebra $\operatorname{\mathcal{H}}_{e}$ excludes the randomness in $\operatorname{\mathcal{T}_{e}^{f}}$. This exclusion is crucial for characterizing the weak dependence structure between epochs.
@@ 4.3 Analysis of the Prototypical Term
Then the prototypical term $\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)$ is exactly $\sum_{e}\operatorname{\mathrm{Bias5}}_{e}$. Our key observation is that, not only
$\mathbb{E}\left[\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left(\mathbb{E}\left[\widehat{\ell}_{t,c}(\pi_{c})\big{|}\operatorname{\mathcal{H}}_{t-1}\right]-\ell_{t,c}(\pi_{c})\right)\right]\leq 0$
@@ Appendix A Useful Lemmas
$L_{e}\triangleq\operatorname{\mathbbm{1}}\left(\max_{c\in[C],a\in[K]}\sum_{t\in\mathcal{T}_{e}}\widetilde{\ell}_{t,c}(a)\leq L+\frac{\iota}{\gamma}\right).$
@@
$\frac{1}{2}\leq\frac{f_{e}(a)+\gamma}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}\leq 2.$
@@
$\widetilde{p}_{t,c}\triangleq\operatorname*{arg\,min}_{p\in\Delta([K])}\left\langle p,\sum_{e^{\prime}=1}^{e-1}\sum_{s\in\mathcal{T}_{e^{\prime}}}\widehat{\ell}_{sc}+\sum_{t^{\prime}\in\mathcal{T}_{e},t^{\prime}<t}\widetilde{\ell}_{t^{\prime}c}\right\rangle-\eta^{-1}F(p)$
@@
It is easy to see that $\widetilde{p}_{t,c}\propto s_{e+1,c}\circ\exp\left(-\eta\sum_{t^{\prime}\in\mathcal{T}_{e},t^{\prime}<t}\widetilde{\ell}_{t^{\prime}c}\right).$
@@
$2s_{e,c}(a)\geq p_{t,c}(a)\geq s_{e,c}(a)/2\qquad\text{and}\qquad 2s_{e,c}(a)\geq\widetilde{p}_{t,c}(a)\geq s_{e,c}(a)/2\,.$
@@
$\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]\leq 4f_{e}(a)\qquad\text{and}\qquad\mathbb{E}_{c\sim\nu}[\widetilde{p}_{t,c}(a)]\leq 4f_{e}(a)\,.$
@@
We define $p_{t}(a)\triangleq\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ and $\widetilde{p}_{t}(a)\triangleq\mathbb{E}_{c\sim\nu}[\widetilde{p}_{t,c}(a)]$ for each time step $t$ and each arm $a$.
@@
$-\frac{\gamma}{f_{e}(a)}\leq\mathbb{E}\left[\frac{f_{e}\left(a\right)-\widehat{f}_{e}\left(a\right)-\frac{1}{2}\gamma}{\widehat{f}_{e}\left(a\right)+\frac{3}{2}\gamma}F_{e}\Big{|}\operatorname{\mathcal{H}}_{e-1}\right]\leq 0.$
@@
$\sum_{t\in\operatorname{\mathcal{T}^{\ell}}}\sum_{c}\Pr(c)\left\langle p_{t,c}-\widetilde{p}_{t,c},\widetilde{\ell}_{t,c}-\widehat{\ell}_{t,c}\right\rangle G\leq\frac{98KT\iota}{L}+\frac{\gamma^{2}LKT}{\iota}.$
@@
Decomposing all time steps into consecutive pairs, specifically, decomposing $\{1,2,\dots,T\}$ into $\{(1,2),(3,4),(5,6),\dots,(t-1,t),\dots,(T-1,T)\}$. Constructing a surrogate loss sequence {~$\{\widetilde{\ell}_{s}\}_{s=1}^{\frac{T}{2}}$ such that for each surrogate time step $s$ the loss vector ~$\widetilde{\ell}_{s}$ is uniformly sampled from the pair of true loss vector $(\ell_{2s-1},\ell_{2s})$. Denote the time step sampled from the pair $(2s-1,2s)$ as $s_{\ell}$. For any constant $\delta\in(0,1)$ and any bandit algorithm such that in each pair of time steps $(t-1,t)$, the algorithm takes actions $a_{t-1}$ and $a_{t}$ from the same distribution $p_{t-1}=p_{t}$, we have
$\sum_{t=1}^{T}\left(\ell_{t,c_{t}}(a_{t})-\ell_{t,c_{t}}(\pi_{c_{t}})\right)\leq 2\sum_{s=1}^{\frac{T}{2}}\left(\widetilde{\ell}_{s,c_{s_{\ell}}}(a_{s_{\ell}})-\widetilde{\ell}_{s,c_{s_{\ell}}}(\pi_{c_{s_{\ell}}})\right)+2\sqrt{T\log(\frac{1}{\delta})}$
@@
$-2\left(\widetilde{\ell}_{s,c_{s_{\ell}}}(a_{s_{\ell}})-\widetilde{\ell}_{s,c_{s_{\ell}}}(\pi_{c_{s_{\ell}}})\right).$
Consider the filtration $\{\widetilde{H}_{s}\}_{s=1}^{\frac{T}{2}}$ such that for each $s$ the $\sigma$-field $\widetilde{H}_{s}$ is generated by the randomness within $c_{t}$ and $a_{t}$ for $t\leq 2s$ and the randomness within sampling from pair $(2\tau-1,2\tau)$ for each $\tau\leq s$. It is easy to see that the sequence$\{Y_{s}\}_{s=1}^{\frac{T}{2}}$ forms a martingale difference sequence adapted to the filtration $\{\widetilde{H}_{s}\}_{s=1}^{\frac{T}{2}}$. Moreover, it is also easy to see that $|Y_{s}|\leq 2$. Using Azuma-Hoeffding’s inequality, for any constant $\delta\in(0,1)$, we have